A study of Navajos residing in the United States formed matched pairs of individuals. In each pair, one member had experienceNo MI MI Diabetes No Diabetes Total Diabetes No Diabetes Total 9 16 25 37 82 119 46 98 144Calculate a point estimate for the odds ratio, that is, for the odds of having a pair with one member MI+Diabetes and the oth

THEN

Calculate a 95% confidence interval for the odds ratio just calcu- laicd.

A study of Navajos residing in the United States formed matched pairs of individuals. In each pair, one member had experienced acute my- ocardial infarction (MI) and the other member was free of heart discase. This rulicd in a toa AA pairs displayd in he able below.
No MI MI Diabetes No Diabetes Total Diabetes No Diabetes Total 9 16 25 37 82 119 46 98 144
Calculate a point estimate for the odds ratio, that is, for the odds of having a pair with one member MI+Diabetes and the other No MI No Diabetes, divided by the odds of having a pair with one member MI+No Diabetes and the other No MI+Diabetes.
Calculate a 95% confidence interval for the odds ratio just calcu- laicd.
Answer #1

Odds ratio =

\frac{9/16}{37/82}=1.24662

The formula for the 95% Confidence Interval for the odds ratio is as follows:

ada-reference.gifДод(OR) t [1 .96xSE( log(OR))] )

The standard error for log(OR) is computed using the following equation:

ada-reference.gifequation_image49.gif

Calculate the natural log of the risk ratio.

log_e(1.24662)=0.22043

Calculate the standard error of the log(OR)

\\SE(log_e(OR)=\sqrt{\frac{1}{a}+\frac{1}{b}+\frac{1}{c}+\frac{1}{c}}\\ =\sqrt{\frac{1}{9}+\frac{1}{16}+\frac{1}{37}+\frac{1}{82}}\\=0.4613385

Calculate the lower and upper confidence bounds on the natural log scale

ada-reference.gif95% conf. İnt.-log (OR)士1.9бы SE(log (OR) )

\\log(95\% conf. int.)=\\=0.2204371897\pm1.96*0.4613385\\=(-0.6837862703,1.12466065)

Convert the log limits back to a normal scale for odds ratios by taking the antilog.

\\e^{-0.6837862703}=0.5047\\ e^{1.12466065}=3.0792\\95\%\ conf.\ int.=(0.5047,3.0792)